OpenStudy (prominer10000):

What is the equation of the line? A. y = -4x B. y = -1/4x C. y = 1/4x D. y = 4x

OpenStudy (will.h):

The slope is obviously negative

OpenStudy (will.h):

you may pick 2 points and use the slope formula to see the slope Slope = y2-y1/x2-x1

OpenStudy (jiteshmeghwal9):

The y-intercept is 0 which implies line is of form y=mx

OpenStudy (will.h):

agreed jit

OpenStudy (prominer10000):

So would it be D?

OpenStudy (will.h):

serious! i gave you a hint and told you the slope is negative

OpenStudy (prominer10000):

Oh sorry, it would be A.

OpenStudy (prominer10000):

Right?

OpenStudy (will.h):

let's see shall we ? pick 2 point (-4,1) (4,-1) slope = y2-y1/x2-x1 slope = -1 - 1/4+ 4 Slope = -2/8 Slope = -1/4

OpenStudy (prominer10000):

Hmm, it must either be B or C..

OpenStudy (will.h):

B

OpenStudy (prominer10000):

Ok, thanks! :)